Skip to content

GATE EC Probability and Statistics questions

Practise Probability and Statistics questions for GATE EC. Explore official PYQs and original topic-wise quizzes, check answers and view available solutions. Free, no sign-in.

Define the sample space and events first. Distinguish independent events from conditional probabilities, and check that the final probability lies between zero and one.

Subject, topic, year and type filter the full library. Match any choice within a filter, and all selected filter categories. Search, source and sort apply to this page.

20 on this page
Official PYQ2026Engineering MathematicsProbability and StatisticsMCQ · 2 marks
Let X,N,YX, N, Y and ZZ be random variables. The variables XX and NN are independent of each other. XX is uniformly distributed between −1-1 and 1;N1 ; N follows Normal distribution with zero mean and unity variance. YY and ZZ are defined as, Y=X+NY=X+N and Z=X2+NZ=X^{2}+N. Which of the following pairs represents the values of correlation …
Official PYQ2025Engineering MathematicsProbability and StatisticsMCQ · 1 mark
A pot contains two red balls and two blue balls. Two balls are drawn from this pot randomly without replacement. What is the probability that the two balls drawn have different colours?
Official PYQ2025Engineering MathematicsProbability and StatisticsMCQ · 2 marks
A source transmits symbol SS that takes values uniformly at random from the set {−2,0,2}\{-2,0,2\}. The receiver obtains Y=S+NY=S+N, where NN is a zero-mean Gaussian random variable independent of SS. The receiver uses the maximum likelihood decoder to estimate the transmitted symbol SS. Suppose the probability of symbol estimation error PeP_{e} is …
Official PYQ2025Engineering MathematicsProbability and StatisticsMSQ · 2 marks
The random variable XX takes values in {−1,0,1}\{-1,0,1\} with probabilities P(X=−1)=P(X=1)=αP(X=-1)=P(X=1)=\alpha and P(X=0)=1−2αP(X=0)=1-2 \alpha, where 0<α<1/20<\alpha<1 / 2. Let g(α)g(\alpha) denote the entropy of XX (in bits), parameterized by α\alpha. Which of the following statements is/are TRUE?
Official PYQ2025Engineering MathematicsProbability and StatisticsNAT · 2 marks
Two fair dice (with faces labeled 1,2,3,4,51,2,3,4,5, and 66) are rolled. Let the random variable XX denote the sum of the outcomes obtained. The expectation of XX is __________ (rounded off to two decimal places).
Official PYQ2024Engineering MathematicsProbability and StatisticsNAT · 1 mark
Suppose X\text{X} and Y\text{Y} are independent and identically distributed random variables that are distributed uniformly in the interval [0,1][0,1]. The probability that X≥Y\text{X} \geq \text{Y} is ______\_\_\_\_\_\_.
Official PYQ2023Engineering MathematicsProbability and StatisticsMCQ · 2 marks
A random variable X\mathrm{X}, distributed normally as N(0,1)\mathrm{N(0,1)} undergoes the transformation Y=h(X)\mathrm{Y}=\mathrm{h}(\mathrm{X}), given in the figure. The form of the probability density function of Y\mathrm{Y} is (In the options given below, a,b,ca, b, c are non-zero constants and g(y)g(y) is piece-wise continuous function).
Official PYQ2022Engineering MathematicsProbability and StatisticsMSQ · 1 mark
Let H(X)H(X) denote the entropy of a discrete random variable XX taking KK possible distinct real values. Which of the following statements is/are necessarily true?
Official PYQ2022Engineering MathematicsProbability and StatisticsMSQ · 2 marks
A state transition diagram with states A,B,A, B, and C,C, and transition probabilities p1,p2,…,p7p_{1}, p_{2}, \dots, p_{7} is shown in the figure (e.g., p1\text{p}_{1} denotes the probability of transition from state AA to BB). For this state diagram, select the statement(s) which is/are universally true.
Official PYQ2021Engineering MathematicsProbability and StatisticsMCQ · 1 mark
Two continuous random variables XX and YY are related as Y=2X+3Y=2X+3 Let σX2\sigma ^{2}_{X} and σY2\sigma ^{2}_{Y}denote the variances of XX and YY, respectively. The variances are related as
Official PYQ2021Engineering MathematicsProbability and StatisticsMCQ · 2 marks
A box contains the following three coins. A fair coin with head on one face and tail on the other face. A coin with heads on both the faces. A coin with tails on both the faces. A coin is picked randomly from the box and tossed. Out of the two remaining coins in the box, one coin is then picked randomly and tossed. If the first toss results in a head, the …
Official PYQ2020Engineering MathematicsProbability and StatisticsNAT · 1 mark
The random variable Y=∫−∞∞W(t)ϕ(t)dt, where ϕ(t)={1;5≤t≤70;otherwiseY=\int_{-\infty }^{\infty }W\left ( t \right )\phi \left ( t \right )dt, \text{ where } \phi \left ( t \right )=\begin{cases} 1; & 5\leq t\leq 7 &\\ 0; & \text{otherwise} \end{cases} and W(t)W(t) is a real white Gaussian noise process with two-sided power spectral density SW(f)=3W/HzS_{W}\left ( f \right )=3 W/Hz, for all ff. …
Official PYQ2020Engineering MathematicsProbability and StatisticsNAT · 1 mark
The two sides of a fair coin are labelled as 00 to 11. The coin is tossed two times independently. Let MM and NN denote the labels corresponding to the outcomes of those tosses. For a random variable XX, defined as X=min(M,N)X = \text{min}(M, N), the expected value E(X)E(X) (rounded off to two decimal places) is ___________.
Official PYQ2020Engineering MathematicsProbability and StatisticsNAT · 2 marks
XX is a random variable with uniform probability density function in the interval [−2, 10][-2,\:10]. For Y=2X−6Y=2X-6, the conditional probability P(Y≤7∣X≥5)P\left ( Y\leq 7\mid X\geq 5 \right ) (rounded off to three decimal places) is __________.
Official PYQ2019Engineering MathematicsProbability and StatisticsNAT · 1 mark
If XX and YY are random variables such that E[2X+Y]=0E\left[2X+Y\right]=0 and E[X+2Y]=33E\left[X+2Y\right]=33, then E[X]+E[Y]=E\left[X\right]+E\left[Y\right]=___________.
Official PYQ2019Engineering MathematicsProbability and StatisticsNAT · 1 mark
Let ZZ be an exponential random variable with mean 11. That is, the cumulative distribution function of ZZ is given by FZ(x)={1−e−xif x≥00if x<0F_{Z}(x)= \left\{\begin{matrix} 1-e^{-x}& \text{if}\: x \geq 0 \\ 0& \text{if}\: x< 0 \end{matrix}\right. Then Pr(Z>2∣Z>1),Pr\left(Z>2 \mid Z>1\right), rounded off to two decimal places, is equal to ___________.
Official PYQ2019Engineering MathematicsProbability and StatisticsNAT · 2 marks
A random variable XX takes values −1-1 and +1+1 with probabilities 0.20.2 and 0.80.8, respectively. It is transmitted across a channel which adds noise N,N, so that the random variable at the channel output is Y=X+NY=X+N. The noise NN is independent of X,X, and is uniformly distributed over the interval [−2,2].[-2,2]. The receiver makes a …
Official PYQ2018Engineering MathematicsProbability and StatisticsNAT · 1 mark
Let X1, X2, X3X_{1},\:X_{2},\:X_{3} and X4X_{4} be independent normal random variable with zero mean and unit variance. The probability that X4X_{4} is the smallest among the four is ________.
Official PYQ2018Engineering MathematicsProbability and StatisticsNAT · 2 marks
A random variable XX takes values −0.5-0.5 and 0.50.5 with probabilities 14\dfrac{1}{4} and 34\dfrac{3}{4}, respectively. The noisy observation of X is Y=X+Z,X\:\text{is}\:Y=X+Z, where ZZ has uniform probability density over the interval (−1,1). X(-1,1).\: X and ZZ are independent. If the MAP rule based detector outputs X^\hat{X} as …
Official PYQ2017Engineering MathematicsProbability and StatisticsNAT · 1 mark
Three fair cubical dice are thrown simultaneously . The probability that all three dice have the same number of dots on the faces showing up is (up to third decimal place)________.